# GraphPilot: Grounded Scene Graph Conditioning for Language-Based Autonomous Driving

### v2

@@ 4.6 Results
Text Actor-Only, SG11 46.4 58.0 0.81 47.1 56.1 0.83 51.7 61.4 0.84 52.2 63.5 0.83
Actor-Only, SG10 46.8 56.2 0.84 51.8 59.4 0.86 51.7 59.8 0.87 51.5 60.0 0.86
JSON Actor-Only, SG11 43.7 55.0 0.80 48.8 55.5 0.87 53.2 62.8 0.84 43.8 51.4 0.88
Actor-Only, SG10 45.7 53.5 0.85 45.5 54.0 0.85 47.3 52.1 0.90 42.2 50.7 0.85
YAML Actor-Only, SG11 45.5 55.7 0.82 44.7 54.6 0.81 52.7 61.6 0.85 52.8 60.4 0.87
Actor-Only, SG10 47.9 57.4 0.83 46.8 54.8 0.84 56.1 64.1 0.87 48.2 55.7 0.85
Mean Actor-Only, SG11 45.2 56.2 0.81 46.9 55.4 0.84 52.5 61.9 0.85 49.6 58.4 0.86
Actor-Only, SG10 46.8 55.7 0.84 48.0 56.1 0.85 51.7 58.7 0.88 47.3 55.5 0.85
Table 5: Extended infraction analysis. Comparison of fine-tuned LMDrive models on safety and behavior sub-metrics, averaged over V3-based configurations. SG10 achieves the best overall balance, reducing collisions and deviations while avoiding the need for test-time scene graphs. Abbreviations: CP = Collision Pedestrians, CV = Collision Vehicles, CL = Collision Layout, RL = Red Light, $S$ = Stop Sign, Off = Off-Road, RD = Route Deviation, TO = Timeout, AB = Agent Blocked. Best results are highlighted in bold.
SG00-PT 42.95 54.81 0.80 0.08 2.83 4.28 2.31 0.00 3.79 11.95 2.93 1.63
SG00-FT 41.55 48.22 0.88 0.00 0.85 2.93 0.10 0.00 2.86 9.94 4.31 2.51
Mean Actor-Only, SG11 47.76 56.78 0.84 0.02 1.35 2.42 1.85 0.02 3.65 5.96 3.63 3.04
Actor-Only, SG10 50.42 57.22 0.87 0.01 1.17 2.07 2.07 0.01 2.65 8.22 2.40 2.55
Scene Graph-Conditioned Fine-Tuning. Tab. 4 reports results for SG11 and SG10 configurations, using the best-performing abstraction (Actor-Only) and injection templates (V2 and V3). For LMDrive, SG10 consistently outperforms SG11 in DS, and V3 surpasses V2, suggesting that exposure to structured context during training enhances internal planning behavior, even when that structure is absent at test-time. The best configuration (Text, Actor-Only, V3, SG10) achieves a DS of 51.8, marking a substantial improvement over the reproduced baseline (42.95) and SG00 fine-tuning (41.55). Moreover, SG10 and SG11 significantly outperform SG01 (44.7), demonstrating that scene graph-conditioned training is more effective than test-time injection only. BEVDriver shows complementary preferences: YAML slightly outperforms Text, and V2 is favored over V3. Nevertheless, SG10 again matches or exceeds SG11, with the top configuration (YAML, Actor-Only, V2, SG10) reaching a DS of 56.1, significantly outperforming both the reproduced baseline (44.70) and SG00 (48.94). We attribute differences in optimal serialization and templates in part to the underlying language backbone (LLaVA-v1.5 in LMDrive vs. LLaMA-7B in BEVDriver), but the broader pattern holds across both: SG10 provides consistent benefits without requiring graph input at test-time.
Taken together, these findings demonstrate that scene graph supervision during training (SG10) not only improves performance but also performs comparably to those using scene graphs at both training and test-time (SG11). This highlights a central insight of our work: models can internalize structured relational information from scene graphs during training and later generalize without requiring such input at test-time. This has significant practical benefits, most notably, SG10 eliminates the need for test-time scene graph generation, which is often complex, error-prone, or costly to deploy in real-time systems. Moreover, removing scene graphs at test-time reduces prompt length, lowering computational load and memory consumption. As a result, SG10 strikes a favorable balance between leveraging structured supervision and maintaining runtime efficiency, making it especially well-suited for real-world autonomous driving applications under resource constraints.
Extended Infraction Analysis. To better understand how scene graph conditioning influences safety and behavior, we analyze detailed sub-metrics in Tab. 5. Rather than focusing on best-case results, we report averages over all V3-based LMDrive configurations (Text, JSON, YAML) to assess the generality of the observed effects.
Compared to the reproduced LMDrive baseline (SG00-PT), SG10 configurations yield significant reductions in critical infractions. Collisions with vehicles and layout elements drop from 2.83 and 4.28 to 1.17 and 2.07, respectively. Collisions with pedestrians are nearly eliminated. Route deviation (RD) also improves notably, dropping from 11.95 to 8.22, indicating better path adherence. These improvements suggest that scene graph injection during training leads to more cautious and structured driving, even without access to graphs at test-time. SG11 achieves similar but slightly worse results than SG10 across most sub-metrics. This could indicate a mild overreliance on explicit graph inputs during deployment or increased prompt complexity. SG00-FT improves some safety-related metrics (e.g., CV and CL), but exhibits the highest RD (9.94), implying weakened command-following precision. Overall, SG10 provides the best balance between route-following accuracy and safety, supporting the idea that scene graph-conditioned training strengthens internal decision making, even when graph inputs are unavailable at test-time.
@@ 5 Conclusion
We presented a model-agnostic approach that grounds language-based autonomous driving models in explicit relational structure through serialized traffic scene graphs. By systematically evaluating different abstraction levels, serialization formats, and injection templates, we identified lightweight and effective strategies that improve driving performance without requiring architectural changes or additional training objectives. Among all configurations, Actor-Only abstractions, which include only dynamic agents and their pairwise interactions, provide a strong balance between semantic relevance and token efficiency. Our experiments further revealed that scene graph prompts enhance planning behavior even when applied only at test-time to pretrained models. More importantly, training with scene graph supervision yields additional consistent performance gains that persist even when scene graphs are removed at test-time. These findings indicate that structural priors can foster more grounded and consistent reasoning in language-based autonomous driving systems.
Limitations. While our approach improves planning through explicit relational supervision, it relies on access to accurate scene graphs during training. Although we show that test-time graph input is not required, generating high-quality scene graphs for training remains a challenge, especially in unstructured or map-less environments. Moreover, we currently inject the serialized scene graph purely at the input level. While this keeps the approach model-agnostic and easy to adopt, future work may explore integrating relational structure more directly within the model backbone to further strengthen visual–relational alignment.
[…]
@@
This supplementary document provides additional information to support the results presented in the main paper. We include detailed implementation settings, full training-time measurements, extended quantitative results, and additional qualitative analyses. All supplementary content is intended to increase reproducibility and offer full transparency of our experimental pipeline.
@@ 6 Implementation Details
This section provides additional details on model configuration, dataset preprocessing, and training hyperparameters that apply to both LMDrive and BEVDriver. These settings remain fixed across all experiments unless explicitly mentioned. Tab. 6 summarizes the shared configuration.
For training and testing, we use eight GPU nodes of a high-performance computing cluster, where each node is equipped with an AMD EPYC 9254 CPU, 384 GB RAM, and an NVIDIA L40s GPU with 48 GB VRAM.
Table 6: Shared Implementation Details for LMDrive and BEVDriver. Both methods use their best-performing language backbones with official pretrained checkpoints (LMDrive: LLaVA-v1.5, BEVDriver: LLaMA-7B). All model, dataset, and training specific hyperparameters are shared across experiments.
@@ 7 Training Time
Serialization Template Full Road-Level Actor-Only
@@
Serialization Template Full Road-Level Actor-Only
@@
Table 9: Computational analysis. Measured latency (in milliseconds) and GPU memory usage (in GB) for each stage of the scene graph pipeline (generation, abstraction, serialization) and model inference across abstraction levels (Full, Road-Level, Actor-Only) and serialization formats (Text, JSON, YAML) under the SG11 system configuration using template V3 and LMDrive as base model. For each measurement, we report the 1st percentile (P01), median (Med), and 99th percentile (P99). The final row (SG10) serves as a baseline where no scene graphs are used at test-time, and thus no scene graph pipeline or additional token overhead is involved.
Serialization Abstraction Generation [ms] Abstraction [ms] Serialization [ms] Inference Time [ms] GPU Usage [GB]
@@
[…]
Across both models, the graph abstraction level is the dominant factor affecting training time, primarily due to differences in sequence length (see Tab. 3 for mean and maximum token counts). The choice of template version has only a negligible impact, and differences between the Text and JSON/YAML serializations follow a consistent pattern: Text yields the shortest training times, while JSON and YAML are comparable but slower due to their higher structural verbosity. The same trends hold for BEVDriver, although its overall training times are substantially shorter than those of LMDrive, which can be attributed to the usage of a different language backbone.
8 Computational Analysis
All experiments in this section were executed on a single NVIDIA L40S GPU node (see Sec. 6). To quantify the computational overhead introduced by scene graph processing, we measure the runtime and resource usage of each component: generation, abstraction, and serialization. These steps are evaluated across all graph abstraction levels (Full, Road-Level, Actor-Only) and serialization formats (Text, JSON, YAML) using template version V3. In addition, since test-time scene graph injection (SG01/SG11) introduces further computational costs, we also report inference latency and GPU memory usage for each configuration with LMDrive as base model.
Tab. 9 presents detailed latency and memory statistics for various SG11 system configurations. Scene graph generation time remains consistent, while abstraction is only applied in Road-Level and Actor-Only configurations, contributing minimally to overall latency. Serialization time varies more significantly: Full and Road-Level abstractions require more tokens and structural elements, especially in JSON and YAML, while Actor-Only remains lightweight.
Inference time is strongly influenced by both abstraction level and serialization format: Text is the most efficient, while JSON and YAML incur higher latency, particularly for Full and Road-Level abstractions due to their verbosity. This trend is also reflected in GPU memory usage. For comparison, we include SG10 system configuration, which excludes scene graph input at test-time, resulting in the lowest overall latency and GPU usage.
9 Detailed Quantitative Results
Extended Baseline Performance. Tab. 10 reports baseline results for LMDrive and BEVDriver without scene graph input (SG00), comparing pretrained (SG00-PT) and fine-tuned (SG00-FT) models across LangAuto Tiny, Short, and Long tracks. This table complements Tab. 1 and provides full benchmark breakdowns for completeness.
Table 11: Extended test-time-only scene graph injection (SG01). Performance across serialization formats, graph abstraction levels, and template versions on different LangAuto benchmarks Tiny, Short, and Long for LMDrive using SG01 system configuration.
Serialization Abstraction Template Tiny Short Long
@@ 9 Detailed Quantitative Results
Extended Test-Time-Only Scene Graph Injection (SG01). Tab. 11 lists complete SG01 results for LMDrive across all serialization formats, abstraction levels, and prompt templates, broken down by LangAuto benchmark track (Tiny, Short, Long). The results complement Tab. 2.
Extended Scene Graph-Conditioned Fine-Tuning. Tab. 12 and Tab. 13 complement Tab. 4 from the main paper by reporting complete benchmark-wise results for LMDrive and BEVDriver, respectively. All configurations use Actor-Only graphs and compare SG11 (train and test-time) and SG01 (test-time-only) settings across serialization formats and prompt templates V2 and V3. While the main analysis focuses on overall means, these detailed results are provided for completeness.
10 Qualitative Results
Alongside this supplementary material, we include a video showcasing qualitative results of our approach, as well as failure cases of the LMDrive and BEVDriver baselines in which our method demonstrates superior driving performance. The failure scenarios include route deviations caused by incorrect instruction following, lane departure due to poor lane tracking, collisions resulting from limited spatial awareness of other traffic participants, and red light violations stemming from failure to recognize traffic signals. Additionally, the video features examples illustrating different abstraction levels of the scene graph (Full, Road-Level, and Actor-Only) for the same driving scenario.
@@ 10 Qualitative Results
Serialization Template Configuration Tiny Short Long
@@
Serialization Template Configuration Tiny Short Long

### v3

@@ 4.6 Results and Ablations
SG00-PT 42.9 54.8 0.80 42.9 54.8 0.80 44.7 49.7 0.90 44.7 49.7 0.90
SG00-FT 41.5 48.2 0.88 41.5 48.2 0.88 48.9 56.4 0.86 48.9 56.4 0.86
Text 46.4 58.0 0.81 47.1 56.1 0.83 51.7 61.4 0.84 52.2 63.5 0.83
JSON 43.7 55.0 0.80 48.8 55.5 0.87 53.2 62.8 0.84 43.8 51.4 0.88
YAML 45.5 55.7 0.82 44.7 54.6 0.81 52.7 61.6 0.85 52.8 60.4 0.87
[…]
Table 5: Noise robustness. Evaluation of SG01 vs. SG11 under varying levels of scene graph corruption, averaged across all serialization formats for LMDrive using template V3.
SG01 SG11
Noise DS RC IS DS RC IS
None 44.7 53.8 0.83 46.9 55.4 0.84
Soft 42.4 52.1 0.82 47.1 55.2 0.84
Medium 42.4 51.4 0.83 46.1 54.7 0.82
Heavy 41.4 51.1 0.82 47.3 56.9 0.83
Table 6: Efficiency analysis. Impact of serialization and abstraction on tokens, latency (ms), and GPU memory (GB) using LMDrive SG11.
Config Tokens Latency GPU
Text Full 623 107 20.9
Road 504 93 19.6
Actor 99 85 18.7
JSON Full 3105 1259 39.5
Road 2242 1066 36.0
Actor 342 351 26.2
YAML Full 1940 1057 32.6
Road 1321 925 29.5
Actor 195 254 25.1
[…]
Explicit Relational Conditioning (SG11). While our primary focus is the internalization of knowledge (SG10), we additionally evaluate the impact of maintaining explicit scene graph conditioning during both training and inference (SG11). As shown in Tab. 4, the benefits of explicit test-time graphs depend on the underlying model architecture and its training protocol. For LMDrive, SG11 generally performs worse than SG10. Since we strictly adhere to the original training recipe where the LLM backbone is kept frozen, the model struggles with the distribution shift introduced by serialized scene graphs during inference, leading to performance interference. In contrast, for BEVDriver, where the LLM is updated via LoRA, SG11 (52.5 mean DS) outperforms SG10 (51.7 mean DS). In this case, the adaptive language backbone better leverages the explicit graph as a constructive reasoning scaffold. These results confirm that while the scene graph provides a valid reasoning signal across models, knowledge internalization (SG10) is a more universally robust and efficient strategy, particularly for architectures with frozen language backbones.
Scene Graph Noise Robustness. We evaluate the reliability of relational grounding by testing SG01 vs. SG11 under three noise levels (soft, medium, heavy) that simulate real-world perception and auto-labeling failures via node and edge dropout (20%–60%) as well as semantic label noise (10%–30%). As shown in Tab. 5, while the zero-shot SG01 configuration exhibits slight degradation as noise increases, SG11 demonstrates exceptional stability, maintaining a consistent 47.3 DS even under heavy noise. This robustness is largely attributed to our use of coarse-grained spatial bins. This discretization provides a layer of symbolic abstraction that is inherently more robust to perception jitter than raw coordinates, as minor fluctuations are absorbed by the quantization buckets. While SG01 is sensitive to label swaps, SG11 learns to treat the scene graph as a relational prior rather than a literal command. During training, scene graph noise acts as a regularizer, forcing the model to cross-reference semantic labels with raw visual features. This learned multi-modal alignment allows SG11 to maintain performance by ”double-checking” noisy labels against visual evidence, making the approach resilient to the corrupted or incomplete graph structures typical of real-world deployment.
Efficiency. We evaluate the computational impact of different scene graph configurations in Sec. 4.6. The results highlight a critical trade-off: while verbose formats such as JSON and YAML provide structured data, they incur substantial latency costs (up to 1259 ms) and GPU memory usage (39.5 GB). In contrast, the Actor-Only Text configuration remains highly efficient. Most importantly, our SG10 strategy matches the baseline (SG00) efficiency perfectly, requiring no additional tokens and maintaining a low latency of 57 ms. This confirms that SG10 provides the benefits of relational grounding with zero added computational overhead during deployment, making it the most viable solution for real-time systems.
@@ 5 Conclusion
We presented a model-agnostic training paradigm that grounds language-based autonomous driving models in explicit relational structure through serialized traffic scene graphs. By systematically evaluating different abstraction levels, serialization formats, and injection templates, we identified lightweight and effective strategies that improve driving performance without requiring architectural changes or additional training objectives. Among all configurations, Actor-Only abstractions, which include only dynamic agents and their pairwise interactions, provide a strong balance between semantic relevance and token efficiency. Our experiments further revealed that scene graph prompts enhance planning behavior, even when applied only at test-time to pretrained models. More importantly, training with scene graph supervision yields additional consistent performance gains that persist even when scene graphs are removed at test-time. These findings indicate that structural priors can foster more grounded and consistent reasoning in language-based autonomous driving systems.
@@
This supplementary material provides additional information to support the results presented in the main paper. We include detailed implementation settings, full training-time measurements, extended quantitative results, and additional qualitative analyses. All supplementary content is intended to enhance reproducibility and provide full transparency into our experimental pipeline.
@@ 6 Implementation Details
This section provides additional details on model configuration, dataset preprocessing, and training hyperparameters for LMDrive, BEVDriver, and SimLingo. We strictly follow the official training strategies and optimization protocols provided by the respective authors for each baseline to ensure a fair and consistent comparison.
Table 7: Implementation Details for LMDrive and BEVDriver. Both methods use their best-performing language backbones with official pretrained checkpoints (LMDrive: LLaVA-v1.5, BEVDriver: LLaMA-7B). We maintain consistent optimization hyperparameters across all experiments, though LoRA parameters apply only to the trainable language backbone of BEVDriver.
@@
LMDrive and BEVDriver share a unified training and optimization infrastructure, as summarized in Table 7. In contrast, SimLingo utilizes a distinct training pipeline and dataset configuration, detailed in Table 8. While all models are trained across eight NVIDIA L40s GPUs, they differ in their trainable parameter subsets: LMDrive keeps the language backbone frozen, whereas BEVDriver and SimLingo utilize LoRA to update the language model. For all experiments, we use a high-performance computing cluster where each node is equipped with an AMD EPYC 9254 CPU, 384 GB RAM, and 48 GB VRAM per GPU.
Table 8: Implementation Details for SimLingo. Optimization strategy for fine-tuning the official pretrained checkpoint using InternVL2-1B as language backbone.
Category Hyperparameter
Model LoRA rank: 32
LoRA alpha: 64
LoRA dropout: 0.3
Dataset data_module: carla_bucket_v12_dreamer
pred_len: 11
use_commentary: True
use_qa: True
qa_augmentation: True
img_shift_augmentation: False
hist_len: 1
route_as: target_point_command
use_lmdrive_commands: True
use_old_towns: True
use_town13: True
use_safety_flag: True
Training Optimizer: AdamW
lr_sched: one_cycle_cosine
lr: 1e-6
warmup_steps: 5% of total steps
weight_decay: 0.2
betas: (0.9, 0.999)
max_epochs: 5
batch_size: 8
gpus: 8
@@ 7 Training Time
Model Serialization Template Full Road-Level Actor-Only
@@
Serialization Template Actor-Only
[…]
Table 11: Computational analysis. Measured latency (in milliseconds) and GPU memory usage (in GB) for each stage of the scene graph pipeline (generation, abstraction, serialization) and model inference across abstraction levels (Full, Road-Level, Actor-Only) and serialization formats (Text, JSON, YAML) under the SG11 system configuration using template V3 and LMDrive as base model. For each measurement, we report the 1st percentile (P01), median (Med), and 99th percentile (P99). The final row (SG10) serves as a baseline in which no scene graphs are used at test time, and thus no scene graph pipeline or additional token overhead is involved.
Serialization Abstraction Generation [ms] Abstraction [ms] Serialization [ms] Latency [ms] GPU [GB]
@@
[…]
Across all models, the graph abstraction level is the dominant factor affecting training time, primarily due to differences in sequence length (see Table 1 in main paper for mean token counts). The choice of the template version has only a negligible impact, and differences between the Text and JSON/YAML serializations follow a consistent pattern: Text yields the shortest training times, while JSON and YAML are comparable but slower due to their higher structural verbosity. Similar trends hold for BEVDriver and SimLingo, although their overall training times are substantially shorter than those of LMDrive. For BEVDriver, this is attributed to the different language backbone, while SimLingo’s efficiency stems from the lightweight InternVL2-1B backbone.
8 Computational Analysis
All experiments in this section were executed on a single NVIDIA L40S GPU node (see Sec. 6). To quantify the computational overhead introduced by scene graph processing, we measure the runtime and resource usage of each component: generation, abstraction, and serialization. These steps are evaluated across all graph abstraction levels (Full, Road-Level, Actor-Only) and serialization formats (Text, JSON, YAML) using template version V3. In addition, since test-time scene graph injection (SG01/SG11) introduces further computational costs, we also report inference latency and GPU memory usage for each configuration with LMDrive as the base model.
Table 11 presents detailed latency and memory statistics for various SG11 system configurations. Scene graph generation time remains consistent, while abstraction is only applied in Road-Level and Actor-Only configurations, contributing minimally to overall latency. Serialization time varies more significantly: Full and Road-Level abstractions require more tokens and structural elements, especially in JSON and YAML, while Actor-Only remains lightweight. Inference time is strongly influenced by both the abstraction level and the serialization format: Text is the most efficient, whereas JSON and YAML incur higher latency, particularly for Full and Road-Level abstractions due to their verbosity. This trend is also reflected in GPU memory usage. For comparison, we include the SG10 system configuration, which excludes scene graph input at test-time, resulting in the lowest overall latency and GPU usage.
9 Detailed Quantitative Results
Extended Baseline Performance. Table 12 reports baseline results for LMDrive and BEVDriver without scene graph input (SG00), comparing pretrained (SG00-PT) and fine-tuned (SG00-FT) models across LangAuto Tiny, Short, and Long tracks. This table complements Table 2 from the main paper and provides full benchmark breakdowns for completeness.
@@ 9 Detailed Quantitative Results
Table 13: Extended zero-shot test-time scene graph injection (SG01). Performance across serialization formats, graph abstraction levels, and template versions on different LangAuto benchmarks Tiny, Short, and Long for LMDrive using SG01 system configuration.
Serialization Graph Abstraction Template Tiny Short Long
@@
Extended Zero-Shot Test-Time Scene Graph Injection (SG01). Table 13 lists complete SG01 results for LMDrive across all serialization formats, abstraction levels, and prompt templates, broken down by LangAuto benchmark track (Tiny, Short, Long). The results complement Table 1 from the main paper.
Extended Scene Graph-Conditioned Fine-Tuning. Table 14 and Table 15 complement Table 2 and  4 from the main paper by reporting complete benchmark-wise results for LMDrive and BEVDriver, respectively. All configurations use Actor-Only graphs and compare SG11 (train and test-time) and SG01 (test-time-only) settings across serialization formats and prompt templates V2 and V3. While the main analysis focuses on overall means, these detailed results are provided for completeness.
@@
Serialization Template Config Tiny Short Long
@@
Serialization Template Config Tiny Short Long
@@
10 Qualitative Results
Alongside this supplementary material, we include a video providing a brief overview of our method and a comparison of the different scene graph abstraction levels (Full, Road-Level, and Actor-Only) for a representative driving scenario. We then showcase qualitative results for each of the three evaluated architectures LMDrive, BEVDriver, and SimLingo. The video highlights specific scenarios where the original pretrained baselines (SG00-PT) fail due to collisions or incorrect instruction following, and demonstrates how our scene graph-conditioned models (SG10) successfully navigate these challenges through internalized relational reasoning.